OpenStudy (anonymous):

equation for point slope form that passes through (-6,0)

OpenStudy (anonymous):

um lets see...point slope is : y-y1=m(x-x1) correct? so the x intercept is obviously -6 and the y wud be 0...idk about the slope though...did u only get one point?

OpenStudy (anonymous):

y-0=m(x+6) is what i have so far